### v1

# DERD-Net: Learning Depth from Event-based Ray Densities

Abstract
Event cameras offer a promising avenue for multi-view stereo depth estimation and Simultaneous Localization And Mapping (SLAM) due to their ability to detect blur-free 3D edges at high-speed and over broad illumination conditions. However, traditional deep learning frameworks designed for conventional cameras struggle with the asynchronous, stream-like nature of event data, as their architectures are optimized for discrete, image-like inputs. We propose a scalable, flexible and adaptable framework for pixel-wise depth estimation with event cameras in both monocular and stereo setups. The 3D scene structure is encoded into disparity space images (DSIs), representing spatial densities of rays obtained by back-projecting events into space via known camera poses. Our neural network processes local subregions of the DSIs combining 3D convolutions and a recurrent structure to recognize valuable patterns for depth prediction. Local processing enables fast inference with full parallelization and ensures constant ultra-low model complexity and memory costs, regardless of camera resolution. Experiments on standard benchmarks (MVSEC and DSEC datasets) demonstrate unprecedented effectiveness: (i) using purely monocular data, our method achieves comparable results to existing stereo methods; (ii) when applied to stereo data, it strongly outperforms all state-of-the-art (SOTA) approaches, reducing the mean absolute error by at least 42%; (iii) our method also allows for increases in depth completeness by more than 3-fold while still yielding a reduction in median absolute error of at least 30%. Given its remarkable performance and effective processing of event-data, our framework holds strong potential to become a standard approach for using deep learning for event-based depth estimation and SLAM. Project page: https://github.com/tub-rip/DERD-Net
1 Introduction
Depth estimation is a fundamental task in computer vision, with key applications in areas such as robotics, autonomous driving, and augmented reality. Traditional stereo vision techniques rely on synchronized cameras to capture images and infer depth by finding correspondences between them. However, these methods often struggle in low-light and fast-motion conditions. Moreover, conventional cameras produce large amounts of redundant data by capturing entire images at fixed intervals, leading to inefficiencies in both data storage and processing.
Unlike conventional cameras, event cameras operate asynchronously, detecting per-pixel brightness changes [22, 8, 9]. This provides high temporal resolution and robustness to motion blur, making them particularly suited for dynamic scenes. Their sparse output enables efficient processing of only relevant areas, making them ideal for real-time tasks such as visual odometry (VO) / SLAM.
Harnessing deep learning for depth estimation with event cameras has the potential to transform such applications. However, adapting neural networks to event data remains challenging due to its asynchronous stream-like nature. Moreover, the scarcity of event camera datasets with ground truth depth [48, 14] results in limited training data, which can lead to overfitting [38]. While simulating data is one way to address this issue, generalizing models trained on simulation to real-world scenarios is far from trivial due to differences in data distribution [27, 36]. Thus, instead of directly processing events (which is prone to overfitting and extensive training time) we rely on intermediate event representations informed by the scene geometry.
One promising approach to capture the 3D scene structure is back-projecting events as rays into space and capturing their intersection densities as disparity space images (DSIs) [31]. DSIs from two or more cameras can be fused, eliminating the need for event synchronization between cameras. This reduces complexity and allows for more robust depth estimation. The Multi-Camera Event-based Multi-View Stereo (MC-EMVS) method has recently produced state-of-the-art (SOTA) results, outperforming other techniques in depth benchmarks across several metrics [13]. To obtain pixel-wise depth estimations from a DSI, MC-EMVS selects the disparity level with the highest ray count, effectively using an argmax operation. A selective threshold filter is applied to predict depth only for pixels with sufficient ray counts.
While straightforward, this approach does not fully utilize the potential of DSIs. It is susceptible to noise and limited in the information it can extract due to ignoring the surrounding pixels and more complex patterns across disparity levels. Consequently, it leads to fewer pixels obtaining depth estimation and less accurate depth predictions than the DSI might potentially allow for. Therefore, there is a need for a more effective approach that can increase the reliability, depth accuracy, and number of pixels for which depth is estimated by recognizing and adequately processing more complex data patterns within the DSIs.
Our Contribution. We propose a novel deep learning framework for event-based depth estimation that is optimized for SLAM scenarios and addresses the aforementioned limitations. An overview is illustrated in Fig. 1. Our approach estimates pixel-wise depth from a DSI using a neural network with 3D convolutions and a recurrent structure. The framework is directly applicable to both monocular and stereo settings. Our key contributions include:
Learning-based Local Processing: For each selected pixel, a small local subregion of the DSI (Sub-DSI) is used as input to the neural network. This novel design leverages the inherent sparsity of event data and DSIs to efficiently process and produce only relevant information for sparse depth-related tasks. (Sec. 3).
Enhanced Data Utilization: Our model captures complex patterns within the DSI, increasing depth prediction accuracy and the number of pixels for which depth can be reliably estimated. Limiting the input to small local subregions around selected pixels enhances generalization by preventing the network from overfitting to specific scenes and augmenting the available training set, as each Sub-DSI serves as an individual data instance.
Efficiency and Scalability: By adopting our Sub-DSI approach, we obtain small independent inputs of fixed size. This enables full parallelization and provides an ultra-light network that has the ability to handle any camera resolution with constant very short inference time. Our network architecture allows the processing of DSIs of variable depth resolution. (Sec. 3.2).
Comprehensive Experiments: We evaluate our model using both monocular and stereo data from standard datasets MVSEC [48] and DSEC [10], employing cross-validation, outperforming the state of the art by a large margin on ten figures of merit. When applied to monocular data, our model achieves performance comparable to SOTA methods that require stereo data (Sec. 4).
To the best of our knowledge, our work is the first learning-based multi-view stereo method to (i) use events and camera poses as input, a crucial requirement for accurate depth prediction over long intervals (as used in SLAM [15]); (ii) demonstrate successful and generalizable depth prediction on real-world event data from DSIs; (iii) report good generalization on all three MVSEC indoor flying sequences [14], even including hybrid methods which combine stereo intensity frames with events.
2 Related Work
Stereo depth estimation with event cameras has been an intriguing problem since the invention of the first event camera by Mahowald and Mead in the 1990s [26, 14, 9] due to their potential for high temporal resolution and robustness to motion blur. Recent approaches have addressed stereo event-based 3D reconstruction for VO and SLAM [45, 47, 46, 7, 15, 35]. These methods assume a static world and known camera motion, using this information to assimilate events over longer time intervals, thereby increasing parallax and producing more accurate semi-dense depth maps. A comprehensive review is provided in [14].
MC-EMVS [13] introduced a novel stereo approach for depth estimation which does not require explicit data association, using DSIs generated from stereo events cameras. By leveraging the sparsity of events and fusing back-projected rays, they outperformed the event-matching based solution of [46] and thereby achieved SOTA results in 3D reconstruction and VO [15]. We advance this approach by employing a compact neural network specialized to derive explicit depth from the DSIs, creating a standardized and effective framework for processing event-based data in deep learning applications that does not rely on event simultaneity or matching.
Deep Learning for depth estimation from event data. Deep learning has significantly advanced depth estimation in traditional monocular and stereo camera setups, achieving remarkable results [21, 20, 37]. However, its application to event camera data remains relatively limited due to the continuous nature of event streams, which require specialized frameworks [14]. For example, in monocular vision, [16] uses synthetic data on a recurrent network to capture temporal information from grid-like event inputs. Yet, mismatches between synthetic and real data degrade performance [32], and monocular depth estimation from events is an ill-posed problem, making high accuracy challenging to achieve with this learning framework [44].
For stereo depth estimation, [38, 40] present two pioneering studies. Specifically, DDES [38] introduced the first deep-learning–based supervised stereo-matching method, while [40] proposed the first unsupervised learning framework. Both methods utilize a First-In First-Out queue to store events at each position, allowing for concurrent time and polarity reservation. Nonetheless, limitations for using visual cues from both left and right event cameras are imposed by the accompanying increase in model complexity and memory costs. Our method overcomes these challenges by maintaining constant input dimensions, defined by the size of the DSI subregion around a selected pixel. It thereby provides a significant advancement in applying deep learning to event-based depth estimation on real-world data.
3 Methodology
In this section, we present our supervised-learning–based approach and the related framework in detail, for which a general overview is provided in Fig. 1. We describe the preprocessing of the data, the architecture of our network (shown in Fig. 2) and the training and inference procedures.
3.1 Framework
As an event camera with pixels moves through a scene, it triggers events and produces a near time-continuous stream of data . Following [31, 13], this stream is sliced by dividing time into intervals. For every time interval, a representative camera viewpoint is selected and a DSI is created. Events are then back-projected into 3D space by casting rays from this viewpoint through the corresponding pixel . The depth axis is discretized into levels, equidistant in inverse linear space, resulting in a 3D DSI of size voxels, whose values represent the number of rays passing through each region (i.e., voxel) of space.
We consider a stereo setup with two synchronized cameras providing two perspectives of the same scene. By leveraging parallax, this configuration enhances depth perception and allows for more accurate 3D reconstruction. For each interval, we construct two DSIs (one for each camera) and fuse them by applying voxel-wise metrics (e.g., harmonic mean) as described in [13]. To compare performance, we also apply our approach to the monocular data of the left camera only.
Since DSIs are typically large and sparse, depth is estimated only for pixels with sufficient information. A confidence map is generated by projecting the DSI onto a 2D grid of size , where each pixel’s value represents the maximum ray density among all depth levels [13] (pixel selection map in Fig. 3). An adaptive Gaussian threshold (AGT) filter is then applied to this grid to select the pixels with a sufficient maximum ray density for reliable depth estimation. For each selected pixel , a surrounding subregion is extracted from the DSI, including the ray counts of all pixels within an L1 radii of :
Each subregion is then normalized individually,
and serves as input to our neural network. Using only small subregions of the DSIs as inputs promotes an efficient and compact architecture operating independent of camera resolution and reduces the risk for overfitting.
The depth estimates are computed in parallel. Since the amount of selected pixels can be controlled by the AGT filter and the dimensions of the Sub-DSIs are fixed, we achieve constant low model complexity and memory costs, regardless of the number of triggered events.
3.2 Network Architecture
The architecture of the neural network is illustrated in Fig. 2, with the dimensions of each layer listed in Tab. 1. The network receives the normalized Sub-DSI (1) as input. The objective is to capture local geometrical patterns in the Sub-DSI to extract more relevant depth information from it than the SOTA argmax approach used in [31, 13]. Since established networks like U-Net [33] often include strong spatial compression and Transformers tend to impose high data demands for reliable generalization [24], we tailor an ultra-lightweight custom architecture that shares design elements with FireNet [34], adapted for very small frame sizes yet variable depth dimensions. Similarly to RAFT-Stereo [23], we adopt convolutions and a Gated Recurrent Unit (GRU) to efficiently handle different depth resolutions, enabling customization of the desired depth precision without modifying architecture.
First, to capture local patterns, further reduce input size and avoid overfitting, we apply a 3D convolutional filter (3D-Conv) with a ReLU activation, a kernel size of and no padding in the spatial dimensions (width and height). For the depth dimension, we set a padding of 1 and a stride of 2 to halve the number of depth layers. We use 4 output channels to capture different patterns simultaneously, resulting in the convolved version of the Sub-DSI (1):
To create the DSIs, rays were cast from the representative camera location into space, passing sequentially through the different depth levels. Since mapping precision requirement may change from scene to scene, we need to deal with variable depth resolution of the DSI. To efficiently and flexibly model this interdependence of consecutive depth layers for a variable , the convolved depth layers are flattened and successively fed into a GRU [6]. The recurrent structure of the GRU allows us to maintain a constant ultra-low count of only 70k parameters in total. It iteratively embeds each depth layer’s information within the context of the previous layers, producing hidden state representations . We then proceed with the final hidden state , which condenses the relevant information from all depth layers along the depth axis:
Finally, a dense layer that preserves the dimension of with ReLU activation and a subsequent output dense layer are applied to process the hidden state. We introduce two versions of the network for custom modification of depth estimation density. In the single-pixel version, the network predicts the depth for the selected pixel while in the multi-pixel version, the output is a grid representing the depth predictions of the central pixel and its 8 neighbors.
Layer Dimensions Details Sub-DSI (Input) Depth Channels Width Height 3D-Convolution K = (3,3,3); P = (1,0,0); S = (2,1,1) ReLU + Flattens Flatten channels and frame GRU Selects final hidden state Dense + ReLU Maintains dimension Dense (Output) or Outputs depth value(s)
3.3 Training and Inference
As supervised loss function for training the neural network model we use the mean absolute error (MAE), with given ground truth depth (this is the case of standard real-world datasets used, such as MVSEC and DSEC – see Sec. 4). To reduce training time and further improve generalization, we additionally employ ensemble learning (EL) [28, 29].
For training, we initialize two identical but independent instances of our neural network with different random weights. The training set is split into two disjoint subsets, enabling parallel training. During testing or inference, each Sub-DSI is processed simultaneously by both networks, and the final depth estimation is obtained by averaging the individual predictions. This helps reduce variance in the predictions, leading to more stable and accurate results. While we use two models, this approach is easily scalable, and incorporating additional models could further improve robustness.
Dataset Dataset Details DSI Parameters Gauss. Filter Training Process Sequences Resolution LiDAR Interval Span Min Dist Max Dist D Sub-DSI Window C Batch Optimizer LR LF Epochs MVSEC Indoor flying 1,2,3 346 260 px 50 ms 1 s 1 m 6.5 m 100 77 99 10 64 AdamW MAE 3 DSEC Zurich04a 640 480 px 100 ms 0.2 s 4 m 50 m 100 77 99 2 64 AdamW MAE 3
4 Experiments
In this section, we evaluate the performance and reliability of the proposed depth estimation approach. Following prior protocols, we conduct experiments on the MVSEC [48] and the DSEC [10] datasets. Ground truth (GT) depth, captured at fixed intervals using LiDAR sensors, serves as reference locations for constructing the respective DSIs over a defined time span. Pixel selection for depth estimation is based on an AGT filter, where the window size determines the surrounding pixel count considered, and a constant is subtracted from the observed ray count.
Supervised training and testing are performed on the network. We investigate the impact of DSIs derived from both monocular and stereo settings during training and testing. Table 2 provides an overview of the key parameters used in the datasets and the training processes of the experiments.
4.1 Metrics
The performance of the networks is evaluated using ten standard metrics commonly employed in depth estimation tasks [13]. We calculate both mean and median errors between the estimated and GT depths, with median errors providing robustness against outliers. Additionally, we report the number of reconstructed points, reflecting the algorithm’s ability to generate valid depth estimations, and the number of outliers (bad-pix [11]), representing the proportion of significant depth estimation errors. In the supplementary, we also compute the scale-invariant logarithmic error (SILog Err) to evaluate the error while considering scale, and the sum of absolute relative differences (AErrR) to assess the relative accuracy of the depth predictions. Finally, we report -accuracy values, which indicate the percentage of points whose estimated depth falls within specified limits relative to GT [41].
4.2 Baseline Methods
We compare our proposed approach against several SOTA methods that have been benchmarked on the task of long-term event-based depth estimation [14], which have been evaluated under the same input conditions (events and camera poses) and output format (semi-dense depth maps) supportive of SLAM.
The Generalized Time-Based Stereovision (GTS) method [18] utilizes a two-step process: first performing stereo matching based on a time-consistency score for each event, followed by depth estimation through triangulation. The Semi-Global Matching (SGM) method [17] is adapted for event-based data by generating time images and subsequently applying stereo matching, with the depth map being refined by masking it at the locations of recent events. Another method, Event-based Stereo Visual Odometry (ESVO) [46], integrates depth estimates by employing Student-t filters, ensuring robust spatio-temporal consistency between stereo time image patches.
The two closest baseline methods for performance comparison of our method are EMVS for monocular vision [31] and MC-EMVS for stereo vision [13]. Both methods extract pixel-wise depth from DSIs by applying the argmax function. To ensure consistency and fairness, we benchmark the methods following the procedure used in prior works [14].
4.3 Experiments on MVSEC Dataset
Algorithm Modality Mean Err Median Err bad-pix #Points [cm] [cm] [%] [million] SOTA EMVS [31] monocular + EMVS [31] monocular + ESVO [46] stereo ESVO indep. 1s stereo SGM indep. 1s stereo GTS indep. 1s stereo MC-EMVS [13] stereo + MC-EMVS [13] stereo + MC-EMVS [13] + MF stereo + Ours DERD-Net monocular + DERD-Net monocular + DERD-Net stereo + DERD-Net stereo + DERD-Net multi-pixel stereo + DERD-Net multi-pixel stereo +
| Scene | Pixel selection map | MC-EMVS [13] | MC-EMVS + Fdenser | DERD-Net (Ours) | Ground truth (GT) |
Algorithm Modality Mean Err Median Err bad-pix #Points [m] [m] [%] [million] SOTA EMVS [31] monocular + EMVS [31] monocular + ESVO [46] stereo MC-EMVS [13] stereo + MC-EMVS [13] stereo + MC-EMVS [13] + MF stereo + Ours DERD-Net monocular + DERD-Net monocular + DERD-Net stereo + DERD-Net stereo + DERD-Net multi-pixel stereo + DERD-Net multi-pixel stereo +
In this section, we describe the experiments conducted on the indoor_flying sequences 1,2,3 of the MVSEC dataset [48] to evaluate the performance of the proposed depth estimation method. We employed three-fold cross-validation by utilizing two sequences for supervised training and reserving the remaining sequence for testing, repeating this process for all three possible combinations of sequences to ensure robustness in our evaluation.
Two pixel-selection filter settings: and . We first trained the single-pixel version of our network on monocular DSIs to compare its performance to EMVS [31]. Subsequently, we retrained it on stereo DSIs fused via the harmonic mean and compared its performance to MC-EMVS [13]. These two baseline methods used an AGT filter with a window of px and a subtractive constant of . Given that our network is designed to extract additional information from the geometrical patterns within the DSI, we hypothesized that it would still produce reliable depth estimates for pixels with lower confidence. To test this hypothesis, we used a larger filter window size of px and a subtractive constant of , resulting in a less strict filter , enabling depth estimation for more pixels. To ensure a representative comparison, we evaluated our networks as well as EMVS and MC-EMVS on both sets of pixels created by and .
Multipixel vs. Morphological Filter. One apparent drawback of MC-EMVS is the limited number of pixels for which depth is estimated compared to other SOTA methods. To address this, [13] presented the option of adding a 4-neighbor morphological filter (MF), which dilates the depth estimation map to increase the number of depth-estimated pixels. We compare this with our framework’s ability to further increase the number of depth-estimated pixels by training and evaluating the multi-pixel version of our network.
Results. The averaged results over all three sequences are displayed in Tab. 3 for all discussed modalities, while the individual results (including performance without EL) are detailed in the Supplementary Material, in Tabs. 9, 10 and 11. Notably, our network’s performance converges after only 3 epochs of training. This rapid convergence is particularly advantageous for future applications where the network might be trained on more heterogeneous datasets or retrained for specific scenarios.
Monocular setup. On the monocular DSIs filtered by , our single-pixel network achieves results comparable to those of stereo SOTA methods and significantly outperforms EMVS [31] by 30% in MAE. Remarkably, even on the 3.27 times larger set of pixels created by , it still achieves better scores than EMVS at for all metrics, except bad-pix. Applying EMVS to the same expanded set of pixels leads to a 76% increase in MAE compared to our framework.
Stereo setup. Applying our single-pixel network to stereo DSIs filtered by allows us to predict depth at significantly more pixels than any other method, except for the SGM method [17], while consistently surpassing all benchmarks across all metrics. The number of pixels increases by 242%, while the MAE and MedAE reduce by 24% and 30%, compared to MC-EMVS [13] with . The only exception is the bad-pix measure, where MC-EMVS performs slightly better. When both methods are compared on the same set of pixels, our approach yields a reduction in both MAE and MedAE of 42% for and 46% for , respectively. Performance remains consistent when using the multi-pixel network, which increases the amount of depth-estimated pixels by a factor of 5.47 for and 4.09 for compared to its single-pixel version, indicating it to be a superior approach to the morphological filter of MC-EMVS, which only rises the number of points by a factor of 3.70 for . As a consequence, the multi-pixel version of our framework estimates depth for almost as many pixels as the SGM method while delivering new SOTA results. Additionally, while no other learning-based methods report good generalization on indoor_flying_2 (see Tab. 10 in the Supplementary Material) due to differences in dynamic characteristics between training and testing events in that split [2, 39], our network excels, underscoring its robust generalization abilities.
Qualitative comparison. To further illustrate the effectiveness of our method, Fig. 3 compares depth maps generated by our single-pixel network against those produced by SOTA method MC-EMVS. Our network not only provides a denser depth estimation, which improves the recognition of contours, but also effectively eliminates the visible outliers produced by MC-EMVS. This improvement is evident when comparing our method to MC-EMVS applied both to the expanded and the original set of pixels, highlighting the robustness and superiority of our approach.
4.4 Experiments on DSEC Dataset
To assess the applicability of our network architecture to different data, we retrained and tested it on DSIs obtained from a stereo setting in the DSEC dataset [10]. This dataset presents unique challenges due to its outdoor driving scenarios, which differ significantly from the indoor environments of the MVSEC dataset, its higher spatial resolution (640 480 px) and different noise characteristics (Prophesee camera vs. DAVIS346 camera). Moreover, straight driving sequences are especially challenging for event-based multi-view stereo due to little motion parallax.
We selected the Zurich_City_04_a sequence to provide a focused in-depth evaluation. We split the sequence into two halves for training and testing. The DSIs were created by fusing the left and right DSIs via the harmonic mean. Analogous to Sec. 4.3, we use the original filter from MC-EMVS [13] with a window size of px and , and a denser filter with a window size of px and . First, the network was trained for 3 epochs on the DSIs of the first half of the sequence and tested on those of the second. The process was then reversed and each network was used to predict in its testing half of the data sequence.
The results of these experiments are displayed in Tab. 4 and illustrated in Fig. 3. Our approach drastically outperforms every other method across all metrics, with our multi-pixel network achieving even slightly better performance than the single-pixel network. For , it reduces the MAE by 55% on a 1.72x higher number of pixels compared to MC-EMVS with a morphological filter. For , depth estimation density is increased by an additional factor of 2.24 while performance remains mostly stable, yielding a reduction in MAE of 62% compared to the argmax operation from MC-EMVS. Remarkably, even on purely monocular DSIs filtered by , our framework achieved superior performance to all benchmarked methods for every metric except MedAE. These results underscore the robustness and versatility of our approach, even in complex real-world outdoor scenes.
Reflection Modality Mean Err Median Err bad-pix #Points [cm] [cm] [%] [million] none stereo + vertical stereo + horizontal stereo + horizontal + vertical stereo + none stereo + vertical stereo + horizontal stereo + horizontal + vertical stereo +
Sub-DSI Modality Mean Err Median Err bad-pix #Points frame size [cm] [cm] [%] [million] stereo + stereo + stereo + stereo + Fdenser stereo + Fdenser stereo + Fdenser
4.5 Data Augmentation Analysis
To further analyze the generalization capabilities of our network, we applied axis-aligned reflections to the DSIs generated from the indoor_flying_1 sequence of the MVSEC dataset. Specifically, we flipped the DSIs horizontally, vertically, and both horizontally and vertically. These transformations allow us to evaluate the network’s performance on scenes that share the same configuration (e.g., distance ranges) but differ due to changes in viewpoint. We retested the single-pixel network that was trained solely on the original indoor_flying 2 and 3 sequences, without retraining it or applying any form of data augmentation during the original training process. This ensures an unbiased evaluation of the network’s inherent ability to generalize to new scenes. Importantly, data augmentation could be used during training to further enhance performance, but for this study, we purposefully avoided it to provide a clear assessment of the network’s natural generalization capabilities.
The results of these experiments are displayed in Tab. 5. Performance only worsened slightly, with results that still significantly outperform all SOTA methods for all tested configurations, indicating that our network might effectively generalize to scenes that share similar depth ranges and texture to those on which it was originally trained.
4.6 Sensitivity Analysis
The hyperparameters used for our experiments are detailed in Tab. 2. The LiDAR interval, the estimated minimum and maximum distances, and the number of depth layers have all been defined to match the protocol in [13]. In Sec. 4.3, we already provided a comparison between the two AGT filters and . In this subsection, we therefore analyze the impact of the size of the Sub-DSI.
We retrained the network for one epoch on the indoor_flying 2 and 3 sequences and compared its test performance on sequence 1 using a radius of 2, 3, and 4 px, effectively creating Sub-DSI frames of size , , and px. From the results reported in Tab. 6 it can be inferred that performance appears to improve as the radius increases. While data size increases proportionally as well, the network was able to achieve better results after a single epoch using a frame size of than when fully trained on frames (see Tab. 9 in the Supplementary Material), highlighting its potential for further performance improvements.
4.7 Runtime
To assess the efficiency of our network, we measured its inference time. It achieved an average of only 0.37 ms per Sub-DSI on an NVIDIA RTX A6000. Since predictions are made independently per pixel, inference for each Sub-DSI can be parallelized on the GPU. The total inference time to estimate a depth map of average density (500 pixels) from MVSEC with is 1.12 ms. This ultra-fast performance, combined with our network’s lightweight architecture, enables efficient execution, making it ideal for real-world applications requiring low-latency depth estimation.
4.8 On robustness compared to other deep-learning–based depth estimation methods
Previous sections have compared our approach to depth estimation methods that have the same input (events and camera poses) and output (semi-dense depth maps), i.e., that tackle the same task (depth prediction for SLAM). This section also provides Tab. 7 to compare to end-to-end learning-based stereo methods, which are “instantaneous” (do not take into account camera poses) and output dense depth. Applying these methods for tracking would therefore require an extra step of depth-pixel selection to avoid becoming computationally expensive.
Since errors come from different observation windows and ground truth points, absolute accuracy is not directly comparable [14]. Nevertheless, comparing the errors in the different splits relative to each other is informative about robustness: we observe that our method is the first one to generalize robustly across all three sequences. No other method reports good generalization on “split 2” of MVSEC because of the difference in dynamic characteristics of events in training and testing on that split (as mentioned in [38, 1]). This is true even when compared to hybrid approaches, despite them also using stereo intensity frames (“2E+2F” input data modality).
Method Modality Split 1 Split 2 Split 3 DDES [38] 2E 16.7 29.4 27.8 EIT-Net [1] 2E 14.2 - 19.4 DTC-SPADE [42] 2E 13.5 - 17.1 Liu et al [25] 2E 20 25 31 StereoSpike [30] 2E 16.5 - 18.4 ASNet [19] 2E 20.46 28.74 22.15 Ghosh et al. [12] 2E 12.1 - 15.6 Chen et al [3] 2E 13.9 - 14.6 StereoFlow-Net [5] 2E 13 - 15 EIS (ICCV 2021) 2E + 2F 13.74 18.43 22.36 SCS-Net [4] 2E + 2F 11.4 - 13.5 N. Uddin et al [40] 2E + 2F 19.7 - 26.4 Zhao et al. [43] 2E + 2F 9.7 - 11.1 DERD-Net 2E 11.69 11.11 12.28
5 Discussion and Conclusion
We have developed the first learning-based multi-view stereo method for event-based depth estimation. Our approach combines input camera poses with events to produce intermediate geometric representations (DSIs) from which depth is estimated using deep learning. It is directly applicable to both monocular and stereo camera setups. By processing small independent subregions of DSIs in parallel, the framework operates independently of camera resolution and facilitates an efficient network under 1 MB in size with an inference time of only 0.37 ms. Its constant low model complexity and memory cost, regardless of the event rate, provide full scalability and enables efficient execution.
Our framework consistently demonstrated superior performance across several metrics compared to other stereo methods and achieved comparable performance when using purely monocular data. It is the first learning-based depth estimation approach that reports robust generalization on all three indoor flying sequences of the MVSEC dataset. Adaptability to different scenes was confirmed on the outdoor driving DSEC dataset, for which it drastically outperformed benchmark approaches across all metrics. Moreover, our framework significantly increased the number of points for which depth can be robustly estimated from DSIs.
Given its exceptional performance, ultra-lightweight architecture, scalability and flexibility across different configurations, our method holds strong potential to become a standard approach for learning depth from events and is highly suitable for real-world applications such as SLAM.
Acknowledgments
This research was funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – EXC 2002/1 “Science of Intelligence” – project number 390523135.
Appendix A Supplementary Material
The averaged results of different SOTA methods compared to DERD-Net are presented in Tab. 8. In Tabs. 9, 10 and 11, the individual performance on each of the respective sequences indoor_flying 1, 2, 3 from the MVSEC dataset are displayed. Table 12 presents the corresponding results for the Zurich_City_04_a sequence from the DSEC dataset. For the data augmentation and sensitivity analysis, the full metrics are disclosed in Tabs. 13 and 14, respectively.
MVSEC Averaged
Algorithm Modality Mean Err Median Err bad-pix SILog Err AErrR log RMSE #Points [cm] [cm] [%] [%] [%] [%] [%] [million] SOTA EMVS [31] monocular + EMVS [31] monocular + ESVO [46] stereo ESVO indep. 1s stereo SGM indep. 1s stereo GTS indep. 1s stereo MC-EMVS [13] stereo + MC-EMVS [13] stereo + MC-EMVS [13] + MF stereo + Ours DERD-Net monocular + DERD-Net monocular + DERD-Net stereo + DERD-Net stereo + DERD-Net multi-pixel stereo + DERD-Net multi-pixel stereo +
MVSEC Indoor Flying 1
Algorithm Modality Mean Err Median Err bad-pix SILog Err AErrR log RMSE #Points [cm] [cm] [%] [%] [%] [%] [%] [million] SOTA EMVS [31] monocular + Forig EMVS [31] monocular + Fdenser ESVO [46] stereo ESVO indep. 1s stereo SGM indep. 1s stereo GTS indep. 1s stereo MC-EMVS [13] stereo + Forig MC-EMVS [13] stereo + Fdenser MC-EMVS [13] + MF stereo + Forig Ours DERD-Net monocular + Forig DERD-Net monocular + Fdenser DERD-Net without EL stereo + Forig DERD-Net stereo + Forig DERD-Net stereo + Fdenser DERD-Net multi-pixel stereo + Forig DERD-Net multi-pixel stereo + Fdenser
MVSEC Indoor Flying 2
Algorithm Modality Mean Err Median Err bad-pix SILog Err AErrR log RMSE #Points [cm] [cm] [%] [%] [%] [%] [%] [million] SOTA EMVS [31] monocular + Forig EMVS [31] monocular + Fdenser ESVO [46] stereo ESVO indep. 1s stereo SGM indep. 1s stereo GTS indep. 1s stereo MC-EMVS [13] stereo + Forig MC-EMVS [13] stereo + Fdenser MC-EMVS [13] + MF stereo + Forig Ours DERD-Net monocular + Forig DERD-Net monocular + Fdenser DERD-Net without EL stereo + Forig DERD-Net stereo + Forig DERD-Net stereo + Fdenser DERD-Net multi-pixel stereo + Forig DERD-Net multi-pixel stereo + Fdenser
MVSEC Indoor Flying 3
Algorithm Modality Mean Err Median Err bad-pix SILog Err AErrR log RMSE #Points [cm] [cm] [%] [%] [%] [%] [%] [million] SOTA EMVS [31] monocular + Forig EMVS [31] monocular + Fdenser ESVO [46] stereo ESVO indep. 1s stereo SGM indep. 1s stereo GTS indep. 1s stereo MC-EMVS [13] stereo + Forig MC-EMVS [13] stereo + Fdenser MC-EMVS [13] + MF stereo + Forig Ours DERD-Net monocular + Forig DERD-Net monocular + Fdenser DERD-Net without EL stereo + Forig DERD-Net stereo + Forig DERD-Net stereo + Fdenser DERD-Net multi-pixel stereo + Forig DERD-Net multi-pixel stereo + Fdenser
DSEC
Algorithm Modality Mean Err Median Err bad-pix SILog Err AErrR log RMSE #Points [m] [m] [%] [%] [%] [%] [%] [million] SOTA EMVS [31] monocular + EMVS [31] monocular + ESVO [46] stereo MC-EMVS [13] stereo + MC-EMVS [13] stereo + MC-EMVS [13] + MF stereo + Ours DERD-Net monocular + DERD-Net monocular + DERD-Net stereo + DERD-Net stereo + DERD-Net multi-pixel stereo + DERD-Net multi-pixel stereo +
Data Augmentation Analysis
Reflection Modality Mean Err Median Err bad-pix SILog Err AErrR log RMSE #Points [cm] [cm] [%] [%] [%] [%] [%] [million] none stereo + vertical stereo + horizontal stereo + horizontal + vertical stereo + none stereo + vertical stereo + horizontal stereo + horizontal + vertical stereo +
Sensitivity Analysis
Sub-DSI Modality Mean Err Median Err bad-pix SILog Err AErrR log RMSE #Points frame size [cm] [cm] [%] [%] [%] [%] [%] [million] stereo + stereo + stereo + stereo + Fdenser stereo + Fdenser stereo + Fdenser

### v2

@@
 Moreover, conventional cameras produce large amounts of redundant data by capturing entire images at fixed intervals, leading to inefficiencies in both data storage and processing.
-Unlike conventional cameras, event cameras operate asynchronously, detecting per-pixel brightness changes [22, 8, 9].
-This provides high temporal resolution and robustness to motion blur, making them particularly suited for dynamic scenes.
+Unlike conventional cameras, event cameras operate asynchronously, detecting per-pixel brightness changes (called “events”) [1, 2, 3].
+This provides high temporal resolution and robustness to motion blur, making them well suited for dynamic scenes.
 Their sparse output enables efficient processing of only relevant areas, making them ideal for real-time tasks such as visual odometry (VO) / SLAM.
@@
 Moreover, the scarcity of event camera datasets with ground truth depth [4, 5] results in limited training data, which can lead to overfitting [6].
-While simulating data is one way to address this issue, generalizing models trained on simulation to real-world scenarios is far from trivial due to differences in data distribution [27, 36].
+While simulating data is one way to address this issue, generalizing models trained on simulation to real-world scenarios is far from trivial due to differences in data distribution [7, 8, 9, 10].
 Thus, instead of directly processing events (which is prone to overfitting and extensive training time) we rely on intermediate event representations informed by the scene geometry.
-One promising approach to capture the 3D scene structure is back-projecting events as rays into space and capturing their intersection densities as disparity space images (DSIs) [31].
+One promising approach for 3D reconstruction is back-projecting events as rays into space and capturing their intersection densities as disparity space images (DSIs) [11].
 DSIs from two or more cameras can be fused, eliminating the need for event synchronization between cameras.
 This reduces complexity and allows for more robust depth estimation.
-The Multi-Camera Event-based Multi-View Stereo (MC-EMVS) method has recently produced state-of-the-art (SOTA) results, outperforming other techniques in depth benchmarks across several metrics [13].
-To obtain pixel-wise depth estimations from a DSI, MC-EMVS selects the disparity level with the highest ray count, effectively using an argmax operation.
+The Multi-Camera Event-based Multi-View Stereo (MC-EMVS) method [12] recently produced state-of-the-art (SOTA) results, outperforming other techniques in depth benchmarks across several metrics.
+To obtain pixel-wise depth estimates from a DSI, MC-EMVS selects the disparity level with the highest ray count, effectively using an argmax operation.
+Ray counting is used as a proxy for finding 3D edges where the rays intersect.
 A selective threshold filter is applied to predict depth only for pixels with sufficient ray counts.
 While straightforward, this approach does not fully utilize the potential of DSIs.
-It is susceptible to noise and limited in the information it can extract due to ignoring the surrounding pixels and more complex patterns across disparity levels.
+It is susceptible to noise and cannot effectively extract cues from surrounding pixels and more complex patterns across disparity levels.
 Consequently, it leads to fewer pixels obtaining depth estimation and less accurate depth predictions than the DSI might potentially allow for.
-Therefore, there is a need for a more effective approach that can increase the reliability, depth accuracy, and number of pixels for which depth is estimated by recognizing and adequately processing more complex data patterns within the DSIs.
+Therefore, we need more effective approaches for extracting depth from DSIs, that are more reliable, accurate, and produce more depth estimates, by adequately recognizing complex ray intersection patterns.
 Our Contribution.
 We propose a novel deep learning framework for event-based depth estimation that is optimized for SLAM scenarios and addresses the aforementioned limitations.
-An overview is illustrated in Fig.
-1.
+An overview is illustrated in Fig.˜1.
 Our approach estimates pixel-wise depth from a DSI using a neural network with 3D convolutions and a recurrent structure.
@@
 This novel design leverages the inherent sparsity of event data and DSIs to efficiently process and produce only relevant information for sparse depth-related tasks.
-(Sec.
-3).
+(Sec.˜3).
 Enhanced Data Utilization: Our model captures complex patterns within the DSI, increasing depth prediction accuracy and the number of pixels for which depth can be reliably estimated.
-Limiting the input to small local subregions around selected pixels enhances generalization by preventing the network from overfitting to specific scenes and augmenting the available training set, as each Sub-DSI serves as an individual data instance.
+Limiting the input to small local subregions around selected pixels enhances generalization by preventing the network from overfitting to dynamics specific to the training scenes and augmenting the available training set, as each Sub-DSI serves as an individual data instance.
 Efficiency and Scalability: By adopting our Sub-DSI approach, we obtain small independent inputs of fixed size.
@@
 Our network architecture allows the processing of DSIs of variable depth resolution.
-(Sec.
-3.2).
-Comprehensive Experiments: We evaluate our model using both monocular and stereo data from standard datasets MVSEC [48] and DSEC [10], employing cross-validation, outperforming the state of the art by a large margin on ten figures of merit.
-When applied to monocular data, our model achieves performance comparable to SOTA methods that require stereo data (Sec.
-4).
-To the best of our knowledge, our work is the first learning-based multi-view stereo method to (i) use events and camera poses as input, a crucial requirement for accurate depth prediction over long intervals (as used in SLAM [15]); (ii) demonstrate successful and generalizable depth prediction on real-world event data from DSIs; (iii) report good generalization on all three MVSEC indoor flying sequences [14], even including hybrid methods which combine stereo intensity frames with events.
+(Sec.˜3.2).
+Comprehensive Experiments: We evaluate our model on both monocular and stereo data from the standard datasets MVSEC [4] and DSEC [13] using cross-validation.
+It outperforms the state of the art by a large margin on ten figures of merit.
+Even using monocular data, our model achieves performance comparable to SOTA methods that require stereo data (Sec.˜4).
+We show downstream applicability and robustness to imperfect (noisy) camera poses.
+To the best of our knowledge, our work is the first learning-based multi-view stereo method to (i) use camera poses along with events as input, which is crucial for accurate depth estimation over long intervals (as used in SLAM [14, 15]); (ii) demonstrate successful and robust depth prediction on real-world event data from DSIs; (iii) report good generalization on all three MVSEC indoor flying sequences [5], even when compared to multi-modal methods that combine stereo intensity frames with events.
+We provide code, trained models and video results for clarity and reproducibility.
 2 Related Work
-Stereo depth estimation with event cameras has been an intriguing problem since the invention of the first event camera by Mahowald and Mead in the 1990s [26, 14, 9] due to their potential for high temporal resolution and robustness to motion blur.
-Recent approaches have addressed stereo event-based 3D reconstruction for VO and SLAM [45, 47, 46, 7, 15, 35].
+Stereo depth estimation with event cameras has been a captivating problem since the invention of the first event camera by Mahowald and Mead in the 1990s [16, 5, 3] due to their potential for high temporal resolution and robustness to motion blur.
+Recent approaches have addressed stereo event-based 3D reconstruction for VO and SLAM [17, 18, 19, 20, 14, 21, 22].
 These methods assume a static world and known camera motion, using this information to assimilate events over longer time intervals, thereby increasing parallax and producing more accurate semi-dense depth maps.
@@
 MC-EMVS [12] introduced a novel stereo approach for depth estimation which does not require explicit data association, using DSIs generated from stereo events cameras.
-By leveraging the sparsity of events and fusing back-projected rays, they outperformed the event-matching based solution of [46] and thereby achieved SOTA results in 3D reconstruction and VO [15].
+By leveraging the sparsity of events and fusing back-projected rays, they outperformed the event-matching–based solution of [19] and thereby achieved SOTA results in 3D reconstruction and VO [14].
+Evidently, this DSI-based 3D reconstruction is robust to imperfect poses estimated using an event camera tracking method.
 We advance this approach by employing a compact neural network specialized to derive explicit depth from the DSIs, creating a standardized and effective framework for processing event-based data in deep learning applications that does not rely on event simultaneity or matching.
@@
 Deep learning has significantly advanced depth estimation in traditional monocular and stereo camera setups, achieving remarkable results [23, 24, 25].
-However, its application to event camera data remains relatively limited due to the continuous nature of event streams, which require specialized frameworks [14].
+However, its application to event camera data remains relatively limited due to the sparse asynchronous nature of event streams, which require specialized frameworks [5].
 For example, in monocular vision, [26] uses synthetic data on a recurrent network to capture temporal information from grid-like event inputs.
-Yet, mismatches between synthetic and real data degrade performance [32], and monocular depth estimation from events is an ill-posed problem, making high accuracy challenging to achieve with this learning framework [44].
+Yet, mismatches between synthetic and real data degrade performance [27], and monocular depth estimation from events is an ill-posed problem, making high accuracy challenging to achieve with this learning-based framework [28].
 For stereo depth estimation, [6, 29] present two pioneering studies.
 Specifically, DDES [6] introduced the first deep-learning–based supervised stereo-matching method, while [29] proposed the first unsupervised learning framework.
-Both methods utilize a First-In First-Out queue to store events at each position, allowing for concurrent time and polarity reservation.
-Nonetheless, limitations for using visual cues from both left and right event cameras are imposed by the accompanying increase in model complexity and memory costs.
-Our method overcomes these challenges by maintaining constant input dimensions, defined by the size of the DSI subregion around a selected pixel.
+Both methods use First-In First-Out queues to store events at each position, allowing for concurrent time and polarity reservation.
+Nevertheless, high event rates lead to greater processing demands and, consequently, increased model complexity and memory requirements, limiting the use of visual cues from both event cameras.
+Our method overcomes these challenges by maintaining constant input dimensions defined by the size of the DSI subregion around a selected pixel, regardless of the event rate.
 It thereby provides a significant advancement in applying deep learning to event-based depth estimation on real-world data.
 3 Methodology
-In this section, we present our supervised-learning–based approach and the related framework in detail, for which a general overview is provided in Fig.
-1.
-We describe the preprocessing of the data, the architecture of our network (shown in Fig.
-2) and the training and inference procedures.
+In this section, we present our supervised-learning–based approach and the related framework in detail, for which a general overview is provided in Fig.˜1.
+We describe the preprocessing of the data, the architecture of our network (shown in Fig.˜2) and the training and inference procedures.
 3.1 Framework
-As an event camera with pixels moves through a scene, it triggers events and produces a near time-continuous stream of data .
-Following [31, 13], this stream is sliced by dividing time into intervals.
-For every time interval, a representative camera viewpoint is selected and a DSI is created.
-Events are then back-projected into 3D space by casting rays from this viewpoint through the corresponding pixel .
-The depth axis is discretized into levels, equidistant in inverse linear space, resulting in a 3D DSI of size voxels, whose values represent the number of rays passing through each region (i.e., voxel) of space.
+As an event camera with pixels moves through a scene, it triggers events and produces a near continuous-time stream of data .
+Following [11, 12], this stream is sliced into time intervals.
+For every time interval, a DSI is created and associated with a camera viewpoint (called Reference Viewpoint) as follows: given camera poses, all events in the interval are back-projected into 3D space by casting rays from the moving camera optical center through the corresponding pixel .
+The depth axis is discretized into levels, equidistant in inverse linear space, resulting in a 3D DSI of size voxels, whose values represent the number of rays passing through each region (i.e., voxel) of space (as shown in Fig.˜1).
+Although input poses are required for building a DSI, they can be obtained from tracking methods or dead-reckoning [30, 12, 14].
 We consider a stereo setup with two synchronized cameras providing two perspectives of the same scene.
@@
 Since DSIs are typically large and sparse, depth is estimated only for pixels with sufficient information.
-A confidence map is generated by projecting the DSI onto a 2D grid of size , where each pixel’s value represents the maximum ray density among all depth levels [13] (pixel selection map in Fig.
-3).
+A confidence map is generated by projecting the DSI onto a 2D grid of size , where each pixel’s value represents the maximum ray density among all depth levels [12] (called pixel selection map in Fig.˜3).
 An adaptive Gaussian threshold (AGT) filter is then applied to this grid to select the pixels with a sufficient maximum ray density for reliable depth estimation.
-For each selected pixel , a surrounding subregion is extracted from the DSI, including the ray counts of all pixels within an L1 radii of :
+For each selected pixel , a surrounding subregion is extracted from the DSI, including the ray counts of all pixels within L1 radii of :
 Each subregion is then normalized individually,
 and serves as input to our neural network.
-Using only small subregions of the DSIs as inputs promotes an efficient and compact architecture operating independent of camera resolution and reduces the risk for overfitting.
+Using only small subregions of the DSIs as inputs leads to an efficient and compact architecture operating independently of camera resolution.
+Furthermore, it reduces the risk of overfitting by encouraging the model to learn generalizable patterns of ray intersections within the Sub-DSIs instead of memorizing semantics and dynamics specific to the training scene.
+Such localized input processing is possible because DSIs consolidate sparse events into a structured format that preserves geometric information within small spatial regions.
 The depth estimates are computed in parallel.
@@
 3.2 Network Architecture
-The architecture of the neural network is illustrated in Fig.
-2, with the dimensions of each layer listed in Tab.
-1.
+Layer Dimensions Details Sub-DSI (Input) Depth Channels Width Height 3D-Convolution K = (3,3,3); P = (1,0,0); S = (2,1,1) ReLU + Flatten Flattens channels and frame GRU Selects final hidden state Dense + ReLU Maintains dimension Dense (Output) or Outputs depth value(s)
+The architecture of the neural network is illustrated in Fig.˜2, with the dimensions of each layer listed in Tab.˜1.
 The network receives the normalized Sub-DSI (1) as input.
-The objective is to capture local geometrical patterns in the Sub-DSI to extract more relevant depth information from it than the SOTA argmax approach used in [31, 13].
+The objective is to capture local geometrical patterns in the Sub-DSI to extract more relevant depth information than the SOTA argmax approach used in [11, 12].
 Since established networks like U-Net [31] often include strong spatial compression and Transformers tend to impose high data demands for reliable generalization [32], we tailor an ultra-lightweight custom architecture that shares design elements with FireNet [33], adapted for very small frame sizes yet variable depth dimensions.
@@
 We introduce two versions of the network for custom modification of depth estimation density.
-In the single-pixel version, the network predicts the depth for the selected pixel while in the multi-pixel version, the output is a grid representing the depth predictions of the central pixel and its 8 neighbors.
-Layer Dimensions Details Sub-DSI (Input) Depth Channels Width Height 3D-Convolution K = (3,3,3); P = (1,0,0); S = (2,1,1) ReLU + Flattens Flatten channels and frame GRU Selects final hidden state Dense + ReLU Maintains dimension Dense (Output) or Outputs depth value(s)
+In the single-pixel version, the network predicts the normalized depth for the selected pixel while in the multi-pixel version, the output is a grid representing the normalized depth predictions of the central pixel and its 8 neighbors.
+Finally, normalized depth is converted into actual depth by mapping to .
 3.3 Training and Inference
-As supervised loss function for training the neural network model we use the mean absolute error (MAE), with given ground truth depth (this is the case of standard real-world datasets used, such as MVSEC and DSEC – see Sec.
-4).
+As supervised loss function for training the neural network model we use the mean absolute error (MAE), with given ground truth depth (this is the case of standard real-world datasets used, such as MVSEC and DSEC – see Sec.˜4).
 To reduce training time and further improve generalization, we additionally employ ensemble learning (EL) [36, 37].
@@
 This helps reduce variance in the predictions, leading to more stable and accurate results.
-While we use two models, this approach is easily scalable, and incorporating additional models could further improve robustness.
+We also present results without EL in Tabs.˜11, 12, 13 and 14 in the Appendix Appendix˜B.
+4 Experiments
 Dataset Dataset Details DSI Parameters Gauss.
-Filter Training Process Sequences Resolution LiDAR Interval Span Min Dist Max Dist D Sub-DSI Window C Batch Optimizer LR LF Epochs MVSEC Indoor flying 1,2,3 346 260 px 50 ms 1 s 1 m 6.5 m 100 77 99 10 64 AdamW MAE 3 DSEC Zurich04a 640 480 px 100 ms 0.2 s 4 m 50 m 100 77 99 2 64 AdamW MAE 3
-4 Experiments
+Filter Training Process Sequences Resolution LiDAR Span D Sub-DSI Window C Batch Optimizer LR LF Epochs MVSEC Indoor flying 346 260 px 50 ms 1 s 1 m 6.5 m 100 77 99 10 64 AdamW MAE 3 DSEC Zurich04a 640 480 px 100 ms 0.2 s 4 m 50 m 100 77 99 2 64 AdamW MAE 3
 In this section, we evaluate the performance and reliability of the proposed depth estimation approach.
@@
 Pixel selection for depth estimation is based on an AGT filter, where the window size determines the surrounding pixel count considered, and a constant is subtracted from the observed ray count.
-Supervised training and testing are performed on the network.
 We investigate the impact of DSIs derived from both monocular and stereo settings during training and testing.
-Table 2 provides an overview of the key parameters used in the datasets and the training processes of the experiments.
+Table˜2 provides an overview of the key parameters used in the datasets and the training processes of the experiments.
+Abbreviations are: minimum depth (), maximum depth (), depth dimensions (), filter window size (Window), subtractive constant (), batch size (Batch), learning rate (LR), and loss function (LF).
 4.1 Metrics
+Method MVSEC DSEC Algorithm Modality Mean Err Median Err bad-pix #Points Mean Err Median Err bad-pix #Points [cm] [cm] [%] [million] [m] [m] [%] [million] SOTA EMVS [11] monocular + EMVS [11] monocular + ESVO [19] stereo SGM [38] stereo GTS [39] stereo MC-EMVS [12] stereo + MC-EMVS [12] stereo + MC-EMVS [12] + MF stereo + Ours DERD-Net monocular + DERD-Net monocular + DERD-Net stereo + DERD-Net stereo + DERD-Net (multi-pixel) stereo + DERD-Net (multi-pixel) stereo +
 The performance of the networks is evaluated using ten standard metrics commonly employed in depth estimation tasks [12].
@@
 Additionally, we report the number of reconstructed points, reflecting the algorithm’s ability to generate valid depth estimations, and the number of outliers (bad-pix [40]), representing the proportion of significant depth estimation errors.
-In the supplementary, we also compute the scale-invariant logarithmic error (SILog Err) to evaluate the error while considering scale, and the sum of absolute relative differences (AErrR) to assess the relative accuracy of the depth predictions.
+In the Appendix, we also compute the scale-invariant logarithmic error (SILog Err) to evaluate the error while considering scale, and the sum of absolute relative differences (AErrR) to assess the relative accuracy of the depth predictions.
 Finally, we report -accuracy values, which indicate the percentage of points whose estimated depth falls within specified limits relative to GT [41].
 4.2 Baseline Methods
-We compare our proposed approach against several SOTA methods that have been benchmarked on the task of long-term event-based depth estimation [14], which have been evaluated under the same input conditions (events and camera poses) and output format (semi-dense depth maps) supportive of SLAM.
+We compare our approach against several SOTA methods that have been benchmarked on the task of long-term event-based depth estimation [5], thus evaluated under the same input conditions (events and camera poses) and output format (semi-dense depth maps) supportive of SLAM.
+In the absence of other deep stereo methods that learn from input camera poses, we also include comparisons against the SOTA instantaneous end-to-end learning-based stereo methods in Sec.˜4.5 for completeness.
+We adopt the same train-test splits established in prior work and standard benchmarks [5].
 The Generalized Time-Based Stereovision (GTS) method [39] utilizes a two-step process: first performing stereo matching based on a time-consistency score for each event, followed by depth estimation through triangulation.
 The Semi-Global Matching (SGM) method [38] is adapted for event-based data by generating time images and subsequently applying stereo matching, with the depth map being refined by masking it at the locations of recent events.
-Another method, Event-based Stereo Visual Odometry (ESVO) [46], integrates depth estimates by employing Student-t filters, ensuring robust spatio-temporal consistency between stereo time image patches.
+Another method, Event-based Stereo Visual Odometry (ESVO) [19] (ESVO2 [42]), integrates depth estimates by employing Student-t filters, ensuring robust spatio-temporal consistency between stereo time image patches.
 The two closest baseline methods for performance comparison of our method are EMVS for monocular vision [11] and MC-EMVS for stereo vision [12].
 Both methods extract pixel-wise depth from DSIs by applying the argmax function.
-To ensure consistency and fairness, we benchmark the methods following the procedure used in prior works [14].
+To ensure consistency and fairness, we benchmark the methods following the procedure established in prior works [5].
 4.3 Experiments on MVSEC Dataset
-Algorithm Modality Mean Err Median Err bad-pix #Points [cm] [cm] [%] [million] SOTA EMVS [31] monocular + EMVS [31] monocular + ESVO [46] stereo ESVO indep.
-1s stereo SGM indep.
-1s stereo GTS indep.
-1s stereo MC-EMVS [13] stereo + MC-EMVS [13] stereo + MC-EMVS [13] + MF stereo + Ours DERD-Net monocular + DERD-Net monocular + DERD-Net stereo + DERD-Net stereo + DERD-Net multi-pixel stereo + DERD-Net multi-pixel stereo +
-| Scene | Pixel selection map | MC-EMVS [13] | MC-EMVS + Fdenser | DERD-Net (Ours) | Ground truth (GT) |
-Algorithm Modality Mean Err Median Err bad-pix #Points [m] [m] [%] [million] SOTA EMVS [31] monocular + EMVS [31] monocular + ESVO [46] stereo MC-EMVS [13] stereo + MC-EMVS [13] stereo + MC-EMVS [13] + MF stereo + Ours DERD-Net monocular + DERD-Net monocular + DERD-Net stereo + DERD-Net stereo + DERD-Net multi-pixel stereo + DERD-Net multi-pixel stereo +
-In this section, we describe the experiments conducted on the indoor_flying sequences 1,2,3 of the MVSEC dataset [48] to evaluate the performance of the proposed depth estimation method.
+| Scene | Pixel selection map | MC-EMVS [12] | MC-EMVS [12] with Fdenser | DERD-Net (Ours) | Ground truth (GT) |
+This section describes the experiments conducted on the indoor_flying sequences 1,2,3 of the MVSEC dataset [4] to evaluate the performance of the proposed depth estimation method.
+Most stereo methods do not evaluate on indoor_flying_4 (because of noisy events from the low-texture floor as the drone flies very low) and the driving sequences (because the stereo baseline is too small for the given depth range and low camera resolution).
 We employed three-fold cross-validation by utilizing two sequences for supervised training and reserving the remaining sequence for testing, repeating this process for all three possible combinations of sequences to ensure robustness in our evaluation.
@@
 Results.
-The averaged results over all three sequences are displayed in Tab.
-3 for all discussed modalities, while the individual results (including performance without EL) are detailed in the Supplementary Material, in Tabs.
-9, 10 and 11.
+The averaged results over all three sequences are displayed in the left half of Tab.˜3 for all discussed modalities, while the individual results (including performance without EL) are detailed in the Appendix, in Tabs.˜12, 13 and 14.
 Notably, our network’s performance converges after only 3 epochs of training.
@@
 As a consequence, the multi-pixel version of our framework estimates depth for almost as many pixels as the SGM method while delivering new SOTA results.
-Additionally, while no other learning-based methods report good generalization on indoor_flying_2 (see Tab.
-10 in the Supplementary Material) due to differences in dynamic characteristics between training and testing events in that split [2, 39], our network excels, underscoring its robust generalization abilities.
 Qualitative comparison.
-To further illustrate the effectiveness of our method, Fig.
-3 compares depth maps generated by our single-pixel network against those produced by SOTA method MC-EMVS.
+To further illustrate the effectiveness of our method, Fig.˜3 compares depth maps generated by our single-pixel network against those produced by SOTA method MC-EMVS.
 Our network not only provides a denser depth estimation, which improves the recognition of contours, but also effectively eliminates the visible outliers produced by MC-EMVS.
@@
 DAVIS346 camera).
-Moreover, straight driving sequences are especially challenging for event-based multi-view stereo due to little motion parallax.
-We selected the Zurich_City_04_a sequence to provide a focused in-depth evaluation.
+Moreover, straight driving sequences are especially challenging for event-based multi-view stereo due to the little motion parallax present in them.
+Setup.
+We select the commonly used Zurich_City_04_a sequence to provide a focused in-depth evaluation.
 We split the sequence into two halves for training and testing.
 The DSIs were created by fusing the left and right DSIs via the harmonic mean.
-Analogous to Sec.
-4.3, we use the original filter from MC-EMVS [13] with a window size of px and , and a denser filter with a window size of px and .
+Analogous to Sec.˜4.3, we use the original filter from MC-EMVS [12] with a window size of px and , and a denser filter with a window size of px and .
 First, the network was trained for 3 epochs on the DSIs of the first half of the sequence and tested on those of the second.
 The process was then reversed and each network was used to predict in its testing half of the data sequence.
-The results of these experiments are displayed in Tab.
-4 and illustrated in Fig.
-3.
+Results.
+The results of these experiments are displayed on the right half of Tab.˜3 and illustrated in Fig.˜3.
 Our approach drastically outperforms every other method across all metrics, with our multi-pixel network achieving even slightly better performance than the single-pixel network.
@@
 These results underscore the robustness and versatility of our approach, even in complex real-world outdoor scenes.
-Reflection Modality Mean Err Median Err bad-pix #Points [cm] [cm] [%] [million] none stereo + vertical stereo + horizontal stereo + horizontal + vertical stereo + none stereo + vertical stereo + horizontal stereo + horizontal + vertical stereo +
-Sub-DSI Modality Mean Err Median Err bad-pix #Points frame size [cm] [cm] [%] [million] stereo + stereo + stereo + stereo + Fdenser stereo + Fdenser stereo + Fdenser
-4.5 Data Augmentation Analysis
-To further analyze the generalization capabilities of our network, we applied axis-aligned reflections to the DSIs generated from the indoor_flying_1 sequence of the MVSEC dataset.
-Specifically, we flipped the DSIs horizontally, vertically, and both horizontally and vertically.
-These transformations allow us to evaluate the network’s performance on scenes that share the same configuration (e.g., distance ranges) but differ due to changes in viewpoint.
-We retested the single-pixel network that was trained solely on the original indoor_flying 2 and 3 sequences, without retraining it or applying any form of data augmentation during the original training process.
-This ensures an unbiased evaluation of the network’s inherent ability to generalize to new scenes.
-Importantly, data augmentation could be used during training to further enhance performance, but for this study, we purposefully avoided it to provide a clear assessment of the network’s natural generalization capabilities.
-The results of these experiments are displayed in Tab.
-5.
-Performance only worsened slightly, with results that still significantly outperform all SOTA methods for all tested configurations, indicating that our network might effectively generalize to scenes that share similar depth ranges and texture to those on which it was originally trained.
-4.6 Sensitivity Analysis
-The hyperparameters used for our experiments are detailed in Tab.
-2.
-The LiDAR interval, the estimated minimum and maximum distances, and the number of depth layers have all been defined to match the protocol in [13].
-In Sec.
-4.3, we already provided a comparison between the two AGT filters and .
-In this subsection, we therefore analyze the impact of the size of the Sub-DSI.
-We retrained the network for one epoch on the indoor_flying 2 and 3 sequences and compared its test performance on sequence 1 using a radius of 2, 3, and 4 px, effectively creating Sub-DSI frames of size , , and px.
-From the results reported in Tab.
-6 it can be inferred that performance appears to improve as the radius increases.
-While data size increases proportionally as well, the network was able to achieve better results after a single epoch using a frame size of than when fully trained on frames (see Tab.
-9 in the Supplementary Material), highlighting its potential for further performance improvements.
-4.7 Runtime
-To assess the efficiency of our network, we measured its inference time.
-It achieved an average of only 0.37 ms per Sub-DSI on an NVIDIA RTX A6000.
-Since predictions are made independently per pixel, inference for each Sub-DSI can be parallelized on the GPU.
-The total inference time to estimate a depth map of average density (500 pixels) from MVSEC with is 1.12 ms.
-This ultra-fast performance, combined with our network’s lightweight architecture, enables efficient execution, making it ideal for real-world applications requiring low-latency depth estimation.
-4.8 On robustness compared to other deep-learning–based depth estimation methods
-Previous sections have compared our approach to depth estimation methods that have the same input (events and camera poses) and output (semi-dense depth maps), i.e., that tackle the same task (depth prediction for SLAM).
-This section also provides Tab.
-7 to compare to end-to-end learning-based stereo methods, which are “instantaneous” (do not take into account camera poses) and output dense depth.
-Applying these methods for tracking would therefore require an extra step of depth-pixel selection to avoid becoming computationally expensive.
-Since errors come from different observation windows and ground truth points, absolute accuracy is not directly comparable [14].
-Nevertheless, comparing the errors in the different splits relative to each other is informative about robustness: we observe that our method is the first one to generalize robustly across all three sequences.
-No other method reports good generalization on “split 2” of MVSEC because of the difference in dynamic characteristics of events in training and testing on that split (as mentioned in [38, 1]).
-This is true even when compared to hybrid approaches, despite them also using stereo intensity frames (“2E+2F” input data modality).
+4.5 Robustness of DERD-Net compared to other deep-learning stereo methods
+Since there are no comparable learning-based methods that use prior camera poses, Tab.˜4 compares end-to-end learning-based stereo methods, which are “instantaneous” (do not take into account camera poses) and output dense depth.
+In order to use their output for efficient VO/SLAM, we would need an extra step of extracting features (keypoints).
+Instead, DERD-Net’s semi-dense depth maps help avoid unnecessary computation by outputting 3D edges for direct visual odometry, as in [14].
 Method Modality Split 1 Split 2 Split 3 DDES [6] 2E 16.7 29.4 27.8 EIT-Net [43] 2E 14.2 - 19.4 DTC-SPADE [44] 2E 13.5 - 17.1 Liu et al [45] 2E 20 25 31 StereoSpike [46] 2E 16.5 - 18.4 ASNet [47] 2E 20.46 28.74 22.15 Ghosh et al.
@@
 [52] 2E + 2F 9.7 - 11.1 DERD-Net 2E 11.69 11.11 12.28
-5 Discussion and Conclusion
+We use the same train-test splits established as the other learning-based methods [5].
+While absolute accuracy is not directly comparable, evaluating the errors in the different splits relative to each other is informative about robustness: we observe that our method is the first one to generalize robustly across all three sequences (Tab.˜4).
+No other method reports good generalization on “split 2” of MVSEC because of the difference in dynamic characteristics of events in training and testing on that split [6, 43].
+This is true even when compared to hybrid approaches, despite them also using stereo intensity frames (“2E+2F” input data modality).
+The observed robustness of our method to such shifts may be supported by the architectural choice of processing only small subregions as input (see Tab.˜2 for Sub-DSI frame size), which encourages the model to learn generalizable patterns within the Sub-DSIs rather than memorizing global scene layout or dataset-specific context.
+4.6 Sensitivity Analyses
+In this section we carry out experiments varying the settings in Tab.˜2. Furthermore, we analyze the robustness of our method to noisy camera poses obtained from an event-based SLAM system.
+Sensitivity with respect to sub-DSI size.
+Varying the horizontal and vertical extent of the Sub-DSIs has an impact on our method’s performance.
+Our experiments show that the performance of DERD-Net can be improved by increasing the frame size of the Sub-DSIs, at the expense of increasing the network complexity (e.g., parameter count and computational cost).
+See Appendix Sec.˜A.1.
+Sensitivity with respect to DSI transformations.
+We analyzed how DERD-Net behaves in the case of previously unseen but structurally similar environments, obtained by means of horizontal and vertical flips of the DSIs.
+Although its performance worsened slightly, it still outperformed all baseline methods.
+This demonstrates robustness to the aforementioned transformations.
+See Appendix Sec.˜A.2.
+Sensitivity with respect to noisy camera poses.
+To assess the importance of having accurate camera poses during DSI construction, we test our framework using noisy poses with drift, mimicking real-world SLAM conditions.
+Instead of ground-truth (GT) poses from LiDAR-IMU odometry, we use poses estimated by the stereo event-based VO system ES-PTAM [14], which reports an Absolute Trajectory Error (ATE) of 131.62 cm over a 50 m-deep scene in the DSEC Zurich_City_04_a sequence.
+Running DERD-Net with these imperfect poses yields the results shown in the top rows of Tab.˜5.
+The percentage values in parentheses denote the relative differences with respect to the performance obtained using ideal (GT) poses (Tab.˜3).
+Remarkably, performance improved across all metrics (likely due to the slight reduction in the number of evaluated points of comparable magnitude), demonstrating strong robustness of DERD-Net to noisy poses obtained from an event-based SLAM system.
+We conduct an additional experiment where the original DERD-Net depth predictions were used to re-estimate the camera poses (in an offline manner, using the camera tracking module in [14]).
+The resulting poses were then used to build DSIs on which DERD-Net was evaluated.
+This “reprojection” loop allows us to assess, using standard depth-based metrics, the robustness of our method to noise in camera poses introduced by DERD-Net’s own depth inaccuracies.
+The results are reported in the bottom rows of Tab.˜5.
+The performance shows only minor degradation, particularly for , with no metric worsening by more than 13%.
+Remarkably, even under such self-induced pose noise, DERD-Net’s depth estimation errors remain roughly 50% lower than those of prior SOTA methods using ideal poses.
+These results demonstrate the practical viability of deploying DERD-Net as a depth-estimation module within a self-sustaining SLAM system.
+Overall, our experiments confirm that DERD-Net remains remarkably robust even when the input poses are significantly degraded, as would be expected in real-world scenarios.
+See also Appendix Secs.˜A.3 and A.4
+Algorithm Poses Filter Mean Err Median Err bad-pix #Points [m] [m] [%] [million] DERD-Net ES-PTAM 1.56 0.45 3.84 1.61 (-3.11%) (-2.17%) (-6.8%) (-3.59%) DERD-Net ES-PTAM 1.74 0.52 4.84 4.49 (-3.33%) (-3.7%) (-3.97%) (-3.23%) DERD-Net Reprojection 1.66 0.49 4.19 1.46 (+3.11%) (+6.52%) (+1.7%) (-12.57%) DERD-Net Reprojection 1.95 0.60 5.65 4.09 (+8.33%) (+11.11%) (+12.1%) (-11.85%)
+4.7 Runtime
+Our network achieved an average inference time of only 0.37 ms per Sub-DSI on an NVIDIA RTX A6000.
+Since predictions are made independently per pixel, inference for each Sub-DSI can be parallelized on the GPU.
+The total inference time to estimate a depth map of average density (500 pixels) from MVSEC with is 1.12 ms.
+Taking MVSEC as an example (DAVIS cameras of 346260 pixels) and DSIs back-projecting 2 million events onto depth planes, then each DSI creation takes 45 ms, DSI fusion takes 26 ms, and pixel selection takes 0.2 ms on an 8-core computer with Intel Xeon(R) W-2225 CPU operating at 4.10 GHz.
+These values are common for both the state-of-the-art method MC-EMVS and DERD-Net.
+It has been shown that DSI creation does not hamper real-time performance [53] because the 3D map can be updated infrequently and on-demand.
+Our network adds only a very small runtime compared to the DSI creation time.
+This ultra-fast performance, combined with its lightweight architecture, enables efficient execution, making DERD-Net ideal for real-world applications requiring low-latency depth estimation.
+5 Conclusion
 We have developed the first learning-based multi-view stereo method for event-based depth estimation.
@@
 By processing small independent subregions of DSIs in parallel, the framework operates independently of camera resolution and facilitates an efficient network under 1 MB in size with an inference time of only 0.37 ms.
-Its constant low model complexity and memory cost, regardless of the event rate, provide full scalability and enables efficient execution.
 Our framework consistently demonstrated superior performance across several metrics compared to other stereo methods and achieved comparable performance when using purely monocular data.
@@
 Moreover, our framework significantly increased the number of points for which depth can be robustly estimated from DSIs.
-Given its exceptional performance, ultra-lightweight architecture, scalability and flexibility across different configurations, our method holds strong potential to become a standard approach for learning depth from events and is highly suitable for real-world applications such as SLAM.
+It also showed strong robustness to noise in camera poses.
+Given its exceptional performance, ultra-lightweight architecture, scalability and flexibility across different configurations, our method holds strong potential to become a standard approach for learning depth from events and is highly suitable for real-world robotic applications requiring low latency and low memory such as SLAM [15].
 Acknowledgments
-This research was funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – EXC 2002/1 “Science of Intelligence” – project number 390523135.
-Appendix A Supplementary Material
-The averaged results of different SOTA methods compared to DERD-Net are presented in Tab.
-8.
-In Tabs.
-9, 10 and 11, the individual performance on each of the respective sequences indoor_flying 1, 2, 3 from the MVSEC dataset are displayed.
-Table 12 presents the corresponding results for the Zurich_City_04_a sequence from the DSEC dataset.
-For the data augmentation and sensitivity analysis, the full metrics are disclosed in Tabs.
-13 and 14, respectively.
+Funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – EXC 2002/1 “Science of Intelligence” – project number 390523135.
+Appendix A Appendix: Sensitivity Analyses
+In this section we report additional experiments to assess our method’s robustness to changes in hyperparameter (Tab.˜6), changes in the scene (Tab.˜7) and noise in camera poses (Tabs.˜8 and 9).
+We also complement the evaluation on the downstream task of camera tracking (Tab.˜10).
+A.1 Sensitivity with respect to Sub-DSI Size
+The hyperparameters used for our experiments are detailed in Tab.˜2.
+The LiDAR’s sampling interval, the estimated minimum and maximum depths, and the number of depth layers have all been defined to match the protocol in [12].
+In Sec.˜4.3, we already provided a comparison between the two AGT filters and .
+In this section, we therefore analyze the impact of the size of the Sub-DSI.
+We retrained the network for one epoch on the indoor_flying 2 and 3 sequences and compared its test performance on sequence 1 using radii of 2, 3, and 4 px, effectively creating Sub-DSI frames of size , , and px, respectively.
+Layer dimensions were adapted, while the overall network architecture remained fixed.
+Sub-DSI Modality Mean Err Median Err bad-pix SILog Err AErrR log RMSE #Points frame size [cm] [cm] [%] [%] [%] [%] [%] [million] stereo + stereo + stereo + stereo + Fdenser stereo + Fdenser stereo + Fdenser
+From the results reported in Tab.˜6 it can be inferred that performance appears to improve as the radii increase.
+The network was able to achieve better results after a single epoch using a frame size of than when fully trained on frames (see Tab.˜12 in this Appendix), highlighting its potential for further performance improvements.
+Nevertheless, increasing the frame size to yielded a reduction of 5% in MAE for both filters after a single training epoch.
+In contrast to that, the network had to apply 65% more 3D-convolutional operations and its total amount of parameters raised from 70k to 270k.
+We therefore decided for a frame size for this study.
+Future research could explore the evident potential to further boost performance by optimizing the sub-DSI size, considering the trade-off between accuracy, parameter count and computational costs.
+A.2 Sensitivity with respect to DSI Transformations
+Next, we analyze the robustness of DERD-Net with respect to transformations of the DSI, in particular axis-aligned reflections of the DSIs generated from the indoor_flying_1 sequence of the MVSEC dataset.
+Specifically, we flipped the DSIs horizontally, vertically, and both horizontally and vertically.
+These transformations effectively generate scenes with similar geometric properties (e.g., distance ranges) but novel spatial configurations.
+This allows us to evaluate how well the network generalizes to previously unseen, yet structurally similar environments.
+We therefore purposely used no data augmentation during training to ensure a representative assessment of the network’s inherent robustness.
+Analogous to previous experiments, we used the single-pixel network that was trained solely on the original indoor_flying 2 and 3 for evaluation.
+No retraining was performed.
+The results of these experiments are displayed in Tab.˜7.
+Performance worsened only slightly, with results that still significantly outperform all SOTA methods for all tested configurations, indicating that our network might effectively generalize to scenes that share similar depth ranges and texture with those on which it was originally trained.
+Reflection Modality Mean Err Median Err bad-pix SILog Err AErrR log RMSE #Points [cm] [cm] [%] [%] [%] [%] [%] [million] none stereo + vertical stereo + horizontal stereo + horizontal + vertical stereo + none stereo + vertical stereo + horizontal stereo + horizontal + vertical stereo +
+A.3 Sensitivity with respect to Noise in Camera Poses
+In Section Sec.˜4.6 we summarized the sensitivity of DERD-Net with respect to noise in the camera poses used to build DSIs, on DSEC data. For completeness, we now show results on MVSEC data.
+We repeat the same experiment as that in the top rows of Tab.˜5 on the MVSEC indoor_flying_1 sequence, for which ES-PTAM reported an ATE of 14.93 cm over a 6 m depth range.
+The results shown in Tab.˜8, and compared to those obtained with GT poses in Tab.˜12, again highlight DERD-Net’s strong robustness to noisy poses estimated by an event-based SLAM system: the MAE and MedAE increased only slightly, while the bad-pix metric even improved.
+The most pronounced decline was in the number of evaluated points.
+Nevertheless, DERD-Net with still predicts depth for 69% more pixels than MC-EMVS with under ideal poses, while achieving a 30% lower MAE.
+For its multi-pixel variant, DERD-Net evaluated with noisy poses maintains superior performance over all state-of-the-art methods using GT poses, while still predicting the highest number of points.
+Algorithm Poses Filter Mean Err Median Err bad-pix #Points [m] [m] [%] [million] DERD-Net ES-PTAM 12.72 6.33 0.65 0.68 (+8.81%) (+16.79%) (-24.42%) (-30.61%) DERD-Net ES-PTAM 15.76 7.36 1.17 1.62 (+6.06%) (+13.76%) (-10.69%) (-46.18%) DERD-Net (multi-pixel) ES-PTAM 13.53 6.76 0.65 3.41 (+10.27%) (+19.01%) (-24.42%) (-33.91%) DERD-Net (multi-pixel) ES-PTAM 16.60 7.65 1.21 6.27 (+6.62%) (+16.08%) (-11.68%) (-46.46%)
+In the interest of thoroughness, we also used poses from the state-of-the-art event-based stereo visual–inertial odometry system ESVO2 [42], which is notably more accurate than ES-PTAM, to run DERD-Net on MVSEC indoor_flying_1, indoor_flying_2, and indoor_flying_3.
+The mean results are reported in Tab.˜9.
+Compared to Tab.˜3, DERD-Net shows only a slight decrease in depth estimation performance on , while it even improves on , confirming its robustness to pose noise from an event-based SLAM system integrating events and inertial data.
+Algorithm Poses Filter Mean Err Median Err bad-pix #Points [m] [m] [%] [million] DERD-Net ESVO2 11.53 5.73 1.09 0.61 (-1.37%) (+4.18%) (+22.47%) (-22.78%) DERD-Net ESVO2 13.69 6.32 1.46 1.45 (-10.17%) (-5.39%) (-14.12%) (-47.65%)
+A.4 Downstream Camera Tracking Performance Analysis
+As intermediate results to those in the bottom part of Tab.˜5, we report offline camera tracking performance using the edge-alignment camera tracking module in [14] acting on input events and the local maps built using DERD-Net’s depth predictions (from GT poses).
+Camera tracking performance is given in terms of ATE and Absolute Rotation Error (ARE) on the DSEC [13] driving dataset in Tab.˜10.
+Although our model was trained only on the Zurich_City_04_a sequence, we evaluate it on all Zurich_City_04 sequences to highlight its generalization capabilities.
+Sequence zc04a zc04b zc04c zc04d zc04e zc04f Duration [s] 35 13.4 53 47.8 13.6 43.1 ATE RMSE [cm] 17.07 7.85 14.00 55.64 5.71 36.11 ARE RMSE [deg] 0.31 0.08 0.45 0.67 0.11 0.72
+The obtained pose errors in the 50 m depth range scenes across all sequences show strong performance of DERD-Net for downstream tasks such as pose estimation via simple photometric edge alignment on event images, as well as its robust generalization even when trained on a single sequence.
+Training on a more diverse set of DSEC sequences would be expected to further enhance these results.
+These values are not comparable to online SLAM tracking results because they assume that the 3D map was pre-built offline using DERD-Net with GT poses.
+Therefore, the estimated camera poses reported here do not accumulate drift.
+Appendix B Appendix - Detailed per-Sequence Results
+The average results of different SOTA methods compared to DERD-Net are presented in Tab.˜11.
+In Tabs.˜12, 13 and 14, the individual performance on each of the respective sequences indoor_flying 1, 2, 3 from the MVSEC dataset are displayed.
+Table˜15 presents the corresponding results for the Zurich_City_04_a sequence from the DSEC dataset.
 MVSEC Averaged
@@
 1s stereo GTS indep.
-1s stereo MC-EMVS [13] stereo + MC-EMVS [13] stereo + MC-EMVS [13] + MF stereo + Ours DERD-Net monocular + DERD-Net monocular + DERD-Net stereo + DERD-Net stereo + DERD-Net multi-pixel stereo + DERD-Net multi-pixel stereo +
+1s stereo MC-EMVS [12] stereo + MC-EMVS [12] stereo + MC-EMVS [12] + MF stereo + Ours DERD-Net monocular + DERD-Net monocular + DERD-Net without EL stereo + DERD-Net stereo + DERD-Net stereo + DERD-Net multi-pixel stereo + DERD-Net multi-pixel stereo +
 MVSEC Indoor Flying 1
@@
 DSEC
-Algorithm Modality Mean Err Median Err bad-pix SILog Err AErrR log RMSE #Points [m] [m] [%] [%] [%] [%] [%] [million] SOTA EMVS [31] monocular + EMVS [31] monocular + ESVO [46] stereo MC-EMVS [13] stereo + MC-EMVS [13] stereo + MC-EMVS [13] + MF stereo + Ours DERD-Net monocular + DERD-Net monocular + DERD-Net stereo + DERD-Net stereo + DERD-Net multi-pixel stereo + DERD-Net multi-pixel stereo +
-Data Augmentation Analysis
-Reflection Modality Mean Err Median Err bad-pix SILog Err AErrR log RMSE #Points [cm] [cm] [%] [%] [%] [%] [%] [million] none stereo + vertical stereo + horizontal stereo + horizontal + vertical stereo + none stereo + vertical stereo + horizontal stereo + horizontal + vertical stereo +
-Sensitivity Analysis
-Sub-DSI Modality Mean Err Median Err bad-pix SILog Err AErrR log RMSE #Points frame size [cm] [cm] [%] [%] [%] [%] [%] [million] stereo + stereo + stereo + stereo + Fdenser stereo + Fdenser stereo + Fdenser
+Algorithm Modality Mean Err Median Err bad-pix SILog Err AErrR log RMSE #Points [m] [m] [%] [%] [%] [%] [%] [million] SOTA EMVS [11] monocular + EMVS [11] monocular + ESVO [19] stereo SGM [38] stereo GTS [39] stereo MC-EMVS [12] stereo + MC-EMVS [12] stereo + MC-EMVS [12] + MF stereo + Ours DERD-Net monocular + DERD-Net monocular + DERD-Net stereo + DERD-Net stereo + DERD-Net multi-pixel stereo + DERD-Net multi-pixel stereo +